OpenStudy (anonymous):

Difference of cubes. x^3 - 2

OpenStudy (anonymous):

What do you want me to do, factor them out?

OpenStudy (anonymous):

\[\mathsf{x^3-a^3=(x-a)(x^2+a^2+xa)}\] \[a = 2^{\frac{1}{3}} \text{ or } \sqrt[3]{2}\]